# Formalizing Computational Paths and Fundamental Groups in Lean

### v1

@@ Abstract
Third, we apply the library to two canonical examples in algebraic topology. We give Lean proofs that the fundamental group of the circle is isomorphic to the integers and that the fundamental group of the torus is isomorphic to the product of two copies of the integers, both via computational paths. These case studies demonstrate that the computational paths approach scales to nontrivial homotopical computations in a modern proof assistant.
@@ 1 Introduction
  • Two full case studies:
@@
    • a computation of the fundamental group of the torus and an isomorphism with the product of two copies of the integers.
@@
Section 2 recalls computational paths and their weak groupoid structure, as well as the basic notions of loop spaces and fundamental groups. Section 3 describes our Lean environment and overall library layout. Section 4 gives the core definitions of paths and rewrites. Section 5 presents the formalization of the weak groupoid structure. Section 6 describes the computation of the fundamental group of the circle. Section 7 does the same for the torus. Section 8 discusses engineering aspects of the development, including library design and automation. Section 9 situates our work in the landscape of homotopy inspired formalizations. Section 10 concludes and sketches directions for future work.
@@ 10 Conclusion and future work
We have presented a Lean 4 formalization of computational paths and their weak groupoid structure. Building on this foundation, we formalized loop spaces and fundamental groups and carried out two classical computations: the fundamental groups of the circle and of the torus. The development is packaged as a reusable library that can serve as a basis for further work on computational paths and homotopy theory in proof assistants.
@@
  • Additional spaces. One can extend the library with further examples such as wedges of circles, higher dimensional tori, and more general cell complexes. This would test the scalability of the computational paths framework and its implementation. In particular, free products of groups arising from wedges of circles are a natural target.
  • Higher structures. The weak groupoid structure considered here is 1 dimensional. Extending the Lean formalization to higher groupoids and higher paths would bring computational paths closer to the higher dimensional structures considered in homotopy type theory. This raises interesting questions about how to represent higher rewrite rules and coherence conditions.
@@
  • Improved automation. The rewrite heavy nature of the development suggests that more sophisticated automation could be helpful. Designing tactics that understand the algebra of paths more deeply is a natural next step. For example, one could implement a tactic that recognizes words in generators and decides equality by normalizing them.

### v2

@@ Abstract
Third, we apply the library to six canonical examples in algebraic topology. We give Lean proofs that the fundamental group of the circle is isomorphic to the integers, the cylinder and Möbius band also have fundamental group isomorphic to the integers (via retraction to the circle), the fundamental group of the torus is isomorphic to the product of two copies of the integers, the fundamental group of the Klein bottle is isomorphic to the semidirect product $\mathbb{Z}\rtimes\mathbb{Z}$, and the fundamental group of the real projective plane is isomorphic to $\mathbb{Z}_{2}$. These case studies demonstrate that the computational paths approach scales to nontrivial homotopical computations in a modern proof assistant.
@@ 1 Introduction
  • Six full case studies:
@@
    • a computation of the fundamental group of the cylinder, showing it retracts to the circle and thus has fundamental group $\mathbb{Z}$,
    • a computation of the fundamental group of the Möbius band, also retracting to the circle with fundamental group $\mathbb{Z}$,
    • a computation of the fundamental group of the torus and an isomorphism with the product of two copies of the integers,
    • a computation of the fundamental group of the Klein bottle and an isomorphism with the semidirect product $\mathbb{Z}\rtimes\mathbb{Z}$,
    • a computation of the fundamental group of the real projective plane and an isomorphism with $\mathbb{Z}_{2}$.
@@
Section 2 recalls computational paths and their weak groupoid structure, as well as the basic notions of loop spaces and fundamental groups. Section 3 describes our Lean environment and overall library layout. Section 4 gives the core definitions of paths and rewrites. Section 5 presents the formalization of the weak groupoid structure. Section 6 describes the computation of the fundamental group of the circle. Section 7 presents the cylinder, which retracts to the circle. Section 8 treats the Möbius band similarly. Section 9 does the same for the torus. Section 10 presents the Klein bottle and its semidirect product fundamental group. Section 11 establishes the fundamental group of the real projective plane. Section 12 discusses engineering aspects of the development, including library design and automation. Section 13 situates our work in the landscape of homotopy inspired formalizations. Section 14 concludes and sketches directions for future work.
@@ 6.2 Isomorphism with the integers
7 Fundamental group of the cylinder
The cylinder $S^{1}\times I$ is homotopy equivalent to the circle, as it deformation retracts onto its central circle [22]. The Lean file ComputationalPaths/Path/HIT/Cylinder.lean [^8] See [2]. models the cylinder as a higher-inductive type with two base points connected by a segment, and loops at each end:
axiom Cylinder : Type u
axiom cylinderBase0 : Cylinder
axiom cylinderBase1 : Cylinder
axiom cylinderSeg : Path cylinderBase0 cylinderBase1
axiom cylinderLoop0 : Path cylinderBase0 cylinderBase0
axiom cylinderLoop1 : Path cylinderBase1 cylinderBase1
The key insight is that the cylinder retracts to either boundary circle. We define a map cylinderToCircle : Cylinder -> Circle that collapses the cylinder to the circle, sending both base points to circleBase and both loops to circleLoop. This map induces an isomorphism on fundamental groups.
7.1 Retraction and encoding
The encoding of cylinder loops proceeds by first mapping to the circle via the retraction, then using the circle’s winding number:
def cylinderLoopToCircleLoop (p : Path cylinderBase0 cylinderBase0) :
Path circleBase circleBase :=
Path.trans (Path.symm (Path.ofEq cylinderToCircle_base0))
(Path.trans (Path.congrArg cylinderToCircle p)
(Path.ofEq cylinderToCircle_base0))
def cylinderEncodePath (p : Path cylinderBase0 cylinderBase0) : Int :=
circleEncodePath (cylinderLoopToCircleLoop p)
The crucial lemma is that cylinderToCircle sends cylinderLoop0 to circleLoop:
theorem cylinderToCircle_loop0 :
cylinderLoopToCircleLoop cylinderLoop0 = circleLoop
This allows all the circle encoding lemmas to be lifted to the cylinder. The decode map constructs loop powers:
def cylinderDecode : Int -> CylinderLoopQuot :=
cylinderLoopZPow
7.2 The equivalence
The encode/decode roundtrip proofs follow the circle pattern:
noncomputable def cylinderPiOneEquivInt :
SimpleEquiv cylinderPiOne Int where
toFun := cylinderEncode
invFun := cylinderDecode
left_inv := cylinderDecode_cylinderEncode
right_inv := cylinderEncode_cylinderDecode
This establishes that $\pi_{1}(\text{Cylinder})\cong\mathbb{Z}$, as expected from the retraction to the circle.
8 Fundamental group of the Möbius band
The Möbius band is a non-orientable surface that, like the cylinder, deformation retracts to its central circle [22]. Despite the twist, its fundamental group is still $\mathbb{Z}$. The Lean file ComputationalPaths/Path/HIT/MobiusBand.lean [^9] See [2]. models it as a higher-inductive type with a single base point and a fundamental loop:
axiom MobiusBand : Type u
axiom mobiusBase : MobiusBand
axiom mobiusLoop : Path mobiusBase mobiusBase
Note that the Möbius band HIT is simpler than the cylinder because we only need the central circle for fundamental group purposes. The twist (non-orientability) would manifest in the transport behavior along the loop, but this does not affect the fundamental group.
8.1 Retraction to the circle
As with the cylinder, we define a retraction mobiusToCircle : MobiusBand -> Circle that sends the base point to circleBase and the central loop to circleLoop:
def mobiusToCircle : MobiusBand -> Circle :=
mobiusBandRec { base := circleBase, loop := circleLoop }
theorem mobiusToCircle_loop :
mobiusLoopToCircleLoop mobiusLoop = circleLoop
The encoding and decoding follow exactly the same pattern as the cylinder:
def mobiusEncodePath (p : Path mobiusBase mobiusBase) : Int :=
circleEncodePath (mobiusLoopToCircleLoop p)
def mobiusDecode : Int -> MobiusLoopQuot :=
mobiusLoopZPow
8.2 The equivalence
The final equivalence establishes the isomorphism:
noncomputable def mobiusPiOneEquivInt :
SimpleEquiv mobiusPiOne Int where
toFun := mobiusEncode
invFun := mobiusDecode
left_inv := mobiusDecode_mobiusEncode
right_inv := mobiusEncode_mobiusDecode
[…]
@@ 9.1 Loop quotients and decode/encode
10 Fundamental group of the Klein bottle
The Klein bottle is a non-orientable surface whose fundamental group is the semidirect product $\mathbb{Z}\rtimes\mathbb{Z}$. Unlike the torus, the two generators do not commute: the characteristic relation is $aba^{-1}=b^{-1}$. The Lean file ComputationalPaths/Path/HIT/KleinBottle.lean [^11] See [2]. encodes this structure as an axiomatic higher-inductive type:
axiom KleinBottle : Type u
axiom kleinBase : KleinBottle
axiom kleinLoopA : Path kleinBase kleinBase
axiom kleinLoopB : Path kleinBase kleinBase
axiom kleinSurf :
Path.trans
(Path.trans kleinLoopA kleinLoopB)
(Path.symm kleinLoopA) =
Path.symm kleinLoopB
The axiom kleinSurf encodes the relation $a\cdot b\cdot a^{-1}=b^{-1}$, which is the defining property of the Klein bottle as opposed to the torus. This relation implies that conjugating the vertical loop $b$ by the horizontal loop $a$ inverts it.
10.1 Code family and transport
The universal cover of the Klein bottle uses the same underlying set as the torus—integer pairs $\mathbb{Z}\times\mathbb{Z}$—but with different transport behavior. The code family kleinCode : KleinBottle -> Type is defined so that:
  • Transport along kleinLoopA sends $(m,n)$ to (m+1,n)(m+1,-n).
  • Transport along kleinLoopB sends $(m,n)$ to $(m,n+1)$.
The negation in the first rule reflects the relation $aba^{-1}=b^{-1}$: going around $a$ and then $b$ is equivalent to going around $b^{-1}$ and then $a$. This is captured by constructing a type equivalence that negates the second coordinate and using the univalence axiom:
def kleinNegateEquiv : SimpleEquiv (Int $\times$$ Int) (Int $\times$$ Int) where
toFun := fun (m, n) => (m, -n)
invFun := fun (m, n) => (m, -n)
left_inv := fun (m, n) => by simp
right_inv := fun (m, n) => by simp
def kleinCodeData : KleinBottleRecData (Type _) where
base := Int $\times$$ Int
loopA := Path.ofEq (ua (kleinShiftNegEquiv)) -- shift and negate
loopB := Path.ofEq (ua (kleinShiftEquiv)) -- shift only
surf := -- proof that the surface relation is satisfied
10.2 Semidirect product structure
The fundamental group $\pi_{1}(K)$ carries the structure of a semidirect product $\mathbb{Z}\rtimes\mathbb{Z}$. The group operation is:
$(m_{1},n_{1})\cdot(m_{2},n_{2})=(m_{1}+m_{2},(-1)^{m_{2}}\cdot n_{1}+n_{2})$
This is implemented in Lean as:
def kleinSemidirectMul (z1 z2 : Int $\times$$ Int) : Int $\times$$ Int :=
(z1.1 + z2.1, Int.negOnePow z2.1 * z1.2 + z2.2)
The key theorem relating loop composition to this multiplication is:
theorem kleinLoopBClass_zpow_mul_loopAClass_zpow (n m : Int) :
LoopQuot.comp (kleinLoopBClass_zpow n) (kleinLoopAClass_zpow m) =
LoopQuot.comp (kleinLoopAClass_zpow m)
(kleinLoopBClass_zpow (Int.negOnePow m * n))
This theorem states that $b^{n}\cdot a^{m}=a^{m}\cdot b^{(-1)^{m}\cdot n}$, which generalizes the surface relation to arbitrary integer powers.
10.3 Encode/decode and the equivalence
The encode map sends a loop to its integer coordinates:
def kleinEncodePath :
Path kleinBase kleinBase -> Int $\times$$ Int :=
fun p => kleinCodeToProd (kleinEncodeRaw kleinBase p)
The decode map constructs a loop from integer coordinates:
def kleinDecodePath (z : Int $\times$$ Int) :
Path kleinBase kleinBase :=
Path.trans (kleinLoopA_zpow z.1) (kleinLoopB_zpow z.2)
The two inverse laws are:
  • kleinEncode_kleinDecodeQuot: $\text{encode}(\text{decode}(m,n))=(m,n)$
  • kleinDecodeQuot_kleinEncode: $\text{decode}(\text{encode}(x))=x$
The final equivalence packages these into a single statement:
noncomputable def kleinPiOneEquivIntProd :
SimpleEquiv kleinPiOne (Int $\times$$ Int) where
toFun := kleinEncode
invFun := kleinDecodeQuot
left_inv := kleinDecodeQuot_kleinEncode
right_inv := kleinEncode_kleinDecodeQuot
This establishes that $\pi_{1}(K)\cong\mathbb{Z}\rtimes\mathbb{Z}$ as sets, with the group structure determined by the semidirect product multiplication.
11 Fundamental group of the real projective plane
The real projective plane $\mathbb{RP}^{2}$ is obtained from a disk by identifying antipodal boundary points. Its fundamental group is the cyclic group of order two, $\mathbb{Z}_{2}$. The Lean file ComputationalPaths/Path/HIT/ProjectivePlane.lean [^12] See [2]. models this as a higher-inductive type with a single loop whose square is trivial:
axiom ProjectivePlane : Type u
axiom projectiveBase : ProjectivePlane
axiom projectiveLoop :
Path projectiveBase projectiveBase
axiom projectiveLoopSquare :
Path.trans projectiveLoop projectiveLoop =
Path.refl projectiveBase
The axiom projectiveLoopSquare states that $\alpha\cdot\alpha=\text{refl}$, where $\alpha$ is the fundamental loop. This is the defining property of $\mathbb{RP}^{2}$: going around the loop twice returns to the identity.
11.1 Code family over Bool
Since the fundamental group has only two elements, the universal cover can be modeled using booleans. The code family projectiveCode : ProjectivePlane -> Type is defined so that:
  • The fiber over the base point is Bool.
  • Transport along projectiveLoop applies boolean negation.
The negation behavior reflects the $\mathbb{Z}_{2}$ structure: going around the loop once switches between the two elements of the cover, and going around twice returns to the starting point.
def projectiveEquiv : SimpleEquiv Bool Bool where
toFun := not
invFun := not
left_inv := Bool.not_not
right_inv := Bool.not_not
theorem projectiveCode_transport_loop (b : Bool) :
Path.transport projectiveLoop
[…]
11.2 Encode and decode
The encode map uses transport in the code family:
def projectiveEncodePath :
Path projectiveBase projectiveBase -> Bool :=
fun p => projectiveCodeToBool
[…]
The key computation rules are:
  • projectiveEncodePath_refl: Encoding the identity path yields false.
  • projectiveEncodePath_loop: Encoding the fundamental loop yields true.
The decode map interprets a boolean as either the identity or the fundamental loop:
def toPathZ2 : Bool -> LoopQuot ProjectivePlane projectiveBase
| false => LoopQuot.id
| true => Quot.mk _ projectiveLoop
11.3 The equivalence
The two inverse laws establish the bijection:
  • projectiveEncode_decode: $\text{encode}(\text{decode}(b))=b$ for $b:\texttt{Bool}$.
  • projectiveDecode_encode: $\text{decode}(\text{encode}(x))=x$ for loops $x$.
The first follows by case analysis on the boolean. The second uses the quotient induction principle and the fact that every loop is rewrite-equivalent to either the identity or the generator.
def projectivePiOneEquivZ2 :
SimpleEquiv (LoopQuot ProjectivePlane projectiveBase) Bool where
toFun := projectiveEncodeQuot
invFun := toPathZ2
left_inv := projectiveDecode_encode
right_inv := projectiveEncode_decode
This completes the proof that $\pi_{1}(\mathbb{RP}^{2})\cong\mathbb{Z}_{2}$, where $\mathbb{Z}_{2}$ is represented as Bool with XOR as addition.
@@ 14 Conclusion and future work
We have presented a Lean 4 formalization of computational paths and their weak groupoid structure. Building on this foundation, we formalized loop spaces and fundamental groups and carried out six classical computations: the fundamental groups of the circle ($\mathbb{Z}$), the cylinder ($\mathbb{Z}$), the Möbius band ($\mathbb{Z}$), the torus ($\mathbb{Z}\times\mathbb{Z}$), the Klein bottle ($\mathbb{Z}\rtimes\mathbb{Z}$), and the real projective plane ($\mathbb{Z}_{2}$). These examples span orientable and non-orientable surfaces, abelian and non-abelian groups, and infinite and finite fundamental groups. The cylinder and Möbius band demonstrate how retraction arguments can be formalized using computational paths. The development is packaged as a reusable library that can serve as a basis for further work on computational paths and homotopy theory in proof assistants.
@@
  • Additional spaces. One can extend the library with further examples such as wedges of circles, higher dimensional tori, spheres of higher dimension, and more general cell complexes. This would test the scalability of the computational paths framework and its implementation. In particular, free products of groups arising from wedges of circles are a natural target, as are the higher homotopy groups $\pi_{n}(S^{n})$.
  • Higher structures. The weak groupoid structure considered here is 1 dimensional. Extending the Lean formalization to higher groupoids and higher paths would bring computational paths closer to the higher dimensional structures considered in homotopy type theory. This raises interesting questions about how to represent higher rewrite rules and coherence conditions. The surface axioms in the Klein bottle and the loop-square axiom in the projective plane are first steps toward handling 2-dimensional path algebra.
@@
  • Improved automation. The rewrite heavy nature of the development suggests that more sophisticated automation could be helpful. Designing tactics that understand the algebra of paths more deeply is a natural next step. For example, one could implement a tactic that recognizes words in generators and decides equality by normalizing them. The Klein bottle development, with its semidirect product structure, particularly benefits from automation that handles the sign-alternating behavior.